Y=-2/3x+3
3x+2y=10

Y=-×-1
Please grapgh these question separately please

Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

Y=-2/3x+3
If you mean y = (-2/3)x + 3
It's a straight line, so you need 2 points.
Pick 2 values for x and find y.
eg,
x = 0, y = 3 --> (0,3)
x = 3, y = 1 --> (3,1)
Plot the 2 points, draw a line thru them.
=================
3x+2y=10
Same as above.
=================
Y=-×-1
Not clear.

Answer by ikleyn(52785)

.
Hello,
if you are a novice in this site, you need/must to know it and to follow it:


    When posting formulas to this forum, use parentheses properly and systematically.    


It is THE RULE of this site.
It is necessary to make your formulas readable and unambiguous.

In other sources, like your textbook etc. the readability is achieved by using special typography means.
This site does not have such means, therefore the only way to make the formulas readable is using parentheses.

I hope that it is sufficient to say it once (one time), and you will follow this requirement.


For example, the way to write your first equation

Y = -2/3x+3

is this:

Y = (-2/3)x+3

(if it is what you want).

And we will know then that it is  NEITHER   Y = -2/(3x+3)   NOR    Y=-2/(3x)+3.